A business firm purchased two machines on 1st January 2017 at a cost of Sh1,500,000 each. Each machine had an estimated useful life of five years and a nil residual value. The straight line method of depreciation is used. On 31st March 2019, one machine was sold for Sh.800,000. The second machine was sold on 1st December 2019 for Sh.250,000.


b) Provision for depreciation account for machinery        

c) Disposal of machinery account                               



1
Expert's answer
2021-04-15T06:57:52-0400

Depreciation per year per machine


"Depreciation=\frac{Value}{number of useful life}"


So, both per year


"Depreciation=\frac{1500000}{5}=300,000"


Total depreciation for 1st machine


"(300000\times2)+(\frac{3}{12}\times300000)=675,000"


Total depreciation for machine 2


"(300000\times2)+(\frac{11}{12}\times300000)=875,000"


Debit depreciation expense 1,550,000

Credit accumulated depreciation 1,550,000

Disposal of machinery

1st machine book value


"1,500,000-675,000=825,000"

2nd machine book value

"1,500,000-875,000=625,000"


On disposal

Debit cash with 1,450,000

Credit machine. 1,450,000

Being disposal of machinery